Session: Jun 2014 (session 2)
Wavelength: 1.3 cm
To give feedback on this experiment, click
(M. Uunila at Wed Jun 18 08:06:17 2014) Success:
(Miguel Sanz at Wed Jun 18 08:07:44 2014) Minor failures: POIN
Since scan nš0284 to nš 0294 stop to redirect antenna.
Scan nš0387 and nš0388 invalid data.
(Carlo Migoni at Wed Jun 18 08:17:05 2014) Success:
No known problems
(M. Uunila at Wed Jun 18 08:27:09 2014) Success:
(K. Borkowski at Wed Jun 18 12:48:38 2014) Minor failures: HIGH
Very high Tsys at times not due to weather, but reason yet unknown. No other known problems.
(Jimmy Cullen at Thu Jun 19 11:36:02 2014) Failure:RECE
Unknown problem, but we suspect a fault in the receiver
(Jun Yang at Mon Jun 23 06:38:50 2014) Success:
It was a blue sky day. No known problems.
(Jun Yang at Mon Jun 23 07:08:45 2014) Success:
Note that atmospheric opacity was corrected in the Tsys data of log and antabfs files at
Onsala. The average opacity along the source direction was 0.26 during the experiment.
(Uwe Bach at Mon Jun 23 14:03:39 2014) Success: